OpenStudy (anonymous):

Consumers in a certain state choose between three long distance telephone services. GTT, NCJ, and Dash. Aggressive marketing by all three companies results in a continual shift of customers among the three services. Each year, GTT loses 25% of its customers to NCJ and 20% to Dash, NCJ loses 5% of its customers to GTT and 20% to Dash, and Dash loses 15% of its customers to GTT and 5 % to NCJ. Assuming that these percentages remain valid over a long period of time, what is each company's expected market share in the long run? Please help me solve this.

OpenStudy (jadebirdfly):

Every year, GTT: 5% to NCJ, 20% to Dash NCJ: 25% to GTT, 25% to Dash Dash: 30% to GTT, 30% to NCJ Now let G, N, D be the proportion of market share of GTT, NCJ, and Dash respectively. We want G = G + 0.25N + 0.3D - 0.05G - 0.2G = 0.75G + 0.25N + 0.3D N = N + 0.05G + 0.3D - 0.25N - 0.25N = 0.5N + 0.05G + 0.3D D = D + 0.2G + 0.25N - 0.3D - 0.3D = 0.4D + 0.2G + 0.25N G + N + D = 1 Simplifying, 0.25G = 0.25N + 0.3D 0.5N = 0.05G + 0.3D 0.6D = 0.2G + 0.25N G + N + D = 1 Intuitively, the 4 equations imply that the net change in market share for each company must be zero. Solving, G = 52.63% N = 21.05% D = 26.32%

OpenStudy (jadebirdfly):

Does this help?

OpenStudy (anonymous):

these numbers are different in my problem. that is the wrong answer

likeabossssssss (likeabossssssss):

@mathmate

likeabossssssss (likeabossssssss):

@Hero @Mehek14

likeabossssssss (likeabossssssss):

@Kainui

OpenStudy (phi):

What course is this? This looks like a markov matrix, and the long-term behavior is given by the eigenvector associated with lambda = 1

OpenStudy (phi):

you should get 20% , 30% , 50% for G, N and D

OpenStudy (anonymous):

@phi that was the answer!!!!! i actually have the same problem with different numbers if you can help me again.... Consumers in a certain state choose between three long distance telephone services. GTT, NCJ, and Dash. Aggressive marketing by all three companies results in a continual shift of customers among the three services. Each year, GTT loses 10% of its customers to NCJ and 25% to Dash, NCJ loses 10% of its customers to GTT and 30% to Dash, and Dash loses 30% of its customers to GTT and 15 % to NCJ. Assuming that these percentages remain valid over a long period of time, what is each company's expected market share in the long run?

OpenStudy (phi):

you should first find the markov matrix, so that \[\left[\begin{matrix}G \\ N \\ D\end{matrix}\right]=A\left[\begin{matrix}G \\ N \\ D\end{matrix}\right]\]

OpenStudy (phi):

can you do that?

OpenStudy (anonymous):

no I'm lost with this. i have asked my professor but haven't heard from him

OpenStudy (phi):

how much does G lose (in total) to its competitors?

OpenStudy (anonymous):

35%?

OpenStudy (phi):

yes, so if it gained nothing from the other 2, how much would it have the next year ?

OpenStudy (phi):

in other words, what percent does it keep?

OpenStudy (anonymous):

65%

OpenStudy (phi):

yes, so it would get G_new= 0.65 G_old

OpenStudy (phi):

\[\left[\begin{matrix}G \\ N \\ D\end{matrix}\right]=\left[\begin{matrix}0.65 & ?& ? \\ ? & ?&?\\ ? & ?&?\end{matrix}\right]\left[\begin{matrix}G \\ N \\ D\end{matrix}\right]\]

OpenStudy (phi):

do you know the G on the left side is calculated by multiplying the top row of the matrix times the vector on the right?

OpenStudy (anonymous):

.65 .10 .30 .10 .60 .15 .25 .30 .55

OpenStudy (anonymous):

i didnt know that. i don't know how to find the vector on the right

OpenStudy (phi):

as a check, the columns should sum to 1 and they do. next, we have to find the eigenvector for eigenvalue 1 do you know how to do that ?

OpenStudy (anonymous):

would it be -.35 .10 .30 .10 -.40 .15 .25 .30 -.45

OpenStudy (phi):

yes, that is a good start A- I (in this case) now find reduced row echelon form

OpenStudy (anonymous):

this is where i got stuck the last time ... how do i find the form

OpenStudy (phi):

step 1: use elimination to zero out the first column can you do that ?

OpenStudy (anonymous):

do i divide the first row by 1/65?

OpenStudy (phi):

-.35 .10 .30 .10 -.40 .15 .25 .30 -.45 you multiply the first row by 0.10/0.35 and add the resultant to the 2nd row

OpenStudy (anonymous):

.025? i added it to the entire row

OpenStudy (phi):

the idea is 0.10/0.35 * -0.35 = -0.10 and when you add that to the 2nd row, the first entry is zeroed out

OpenStudy (phi):

0.10/0.35 is 0.28571 when you multiply the first row by that you get -0.100000 0.028571 0.085714 when you add that row to the 2nd row you get 0.00000 -0.37143 0.23571

OpenStudy (anonymous):

ok i understand now.... what is the next step

OpenStudy (phi):

you use the same idea to zero the first entry in the last row

OpenStudy (anonymous):

how did you get -0.37143 0.23571 in the second row?

OpenStudy (phi):

starting at the beginning multiply the first row by 0.10/0.35 you get -0.100000 0.028571 0.085714

OpenStudy (phi):

you then add that to the 2nd row -0.100000 0.028571 0.085714 .10 -.40 .15

OpenStudy (anonymous):

ok then i know you add -0.100000 to the first one in the second row giving you 0

OpenStudy (phi):

you get a new 2nd row 0.00000 -0.37143 0.23571

OpenStudy (anonymous):

ok .. i will try the 3rd row now

OpenStudy (anonymous):

-0.100000 0.028571 0.085714 0.00000 -0.37143 0.23571 .15 0.328571 -.364286

OpenStudy (anonymous):

is that correct?

OpenStudy (phi):

you leave the first row as it was. The -0.100000 0.028571 0.085714 is only a "side calculation" in other words, after we zero out the first entry in the 2nd row *only* the 2nd row in the matrix changes. when you do the 3rd row, you still use the *original* first row.

OpenStudy (phi):

here is what you have after doing the first row, but before doing the 3rd row -.35 .10 .30 0.00000 -0.37143 0.23571 .25 .30 -.45

OpenStudy (phi):

*after doing the 2nd row, but before doing the 3rd row

OpenStudy (anonymous):

yes i forgot you multiply ... i was adding what is the next step

OpenStudy (phi):

zero out the 0.25 in the 3rd row you do that by multiplying the first row by 0.25/0.35 to get a *temporary row" that you add to the 3rd row to get a *new third row*

OpenStudy (anonymous):

is the third row .... 0 .37142857 -0.23571429

OpenStudy (phi):

yes. good news, it looks close to the 2nd row

OpenStudy (anonymous):

ok what is the second row

OpenStudy (anonymous):

what is the next step***

OpenStudy (phi):

-.35 .10 .30 0.00000 -0.37143 0.23571 .25 .30 -.45 becomes -.35 .10 .30 0.00000 -0.37143 0.23571 0 .37142857 -0.23571429

OpenStudy (anonymous):

yes i have that

OpenStudy (phi):

we go to the next column over, and down 1 row the new "pivot" is the -0.37143 multiply the 2nd row by 1 (i.e. do nothing) and add it to the 3rd row. This will eliminate the entry below the pivot

OpenStudy (phi):

you will get -.35 .10 .30 0.00000 -0.37143 0.23571 0.000 0.00000 0.0000

OpenStudy (phi):

next, we want the pivots to be 1 the first pivot is the -0.35 in location (1,1) to make it one, we divide the 1st row (all entries) by -0.35 to get a *new first row*

OpenStudy (anonymous):

1 .28571427 .85714286 0.00000 -0.37143 0.23571 0.000 0.00000 0.0000

OpenStudy (anonymous):

is that correct?

OpenStudy (phi):

except for the signs. they should be negative for the other 2 entries, right ?

OpenStudy (anonymous):

correct i was diving by .35 not -.35

OpenStudy (anonymous):

dividing*

OpenStudy (phi):

1 - 0.28571427 - 0.85714286 0.00000 -0.37143 0.23571 0.000 0.00000 0.0000 now make the 2nd pivot 1. that is, divide the 2nd row by -0.37143 to get a new 2nd row

OpenStudy (anonymous):

1 - 0.28571427 - 0.85714286 -0.37143 1 -0.63460141 0.000 0.00000 0.0000

OpenStudy (phi):

the first entry in the 2nd row stays 0

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

1 - 0.28571427 - 0.85714286 0 1 -0.63460141 0.000 0.00000 0.0000

OpenStudy (phi):

finally, we want the pivot columns to be all zeros (except for the pivot) in other words, we want to "clear out" the -0.28... above the 2nd pivot to do that, multiply the 2nd row by 0.2857... and add that to the first row to get a *new 1st row* (the 2nd row stays as is)

OpenStudy (anonymous):

0.2857 0 0.67582818 0 1 -0.63460141 0.000 0.00000 0.0000

OpenStudy (phi):

you did something wrong. the 2nd row (temp result) should be 0 0.28571427 -0.1813... now add that to the 1st row 1 - 0.28571427 - 0.85714286

OpenStudy (anonymous):

oh ok

OpenStudy (phi):

what is the matrix ?

OpenStudy (anonymous):

1 0 - 1.03544286 -0.37143 1 -0.63460141 0.000 0.00000 0.0000

OpenStudy (anonymous):

sorry this.... 1 0 - 1.03544286 0 1 -0.63460141 0.000 0.00000 0.0000

OpenStudy (phi):

better!

OpenStudy (phi):

now we "read off" the eigenvector or we solve for the eigenvector

OpenStudy (anonymous):

how?

OpenStudy (phi):

do you know how to find vectors in the "null space" of a matrix? (or have you heard of this?)

OpenStudy (anonymous):

sorry i have not

OpenStudy (phi):

not to get to far off track: the eigenvector equation is \[ A x = \lambda x \\ Ax - \lambda x = 0 \\ \left( A - \lambda I\right) x = 0 \] and x is a vector that gives 0 (i.e. a vector of zeros) when we multiply by the matrix \( \left( A - \lambda I\right) \)

OpenStudy (phi):

all of which means reduced row echelon form matrix we found times the the eigenvector x gives 0 \[\left[\begin{matrix}1 & 0 & a \\ 0 & 1 & b \\ 0 & 0 & 0\end{matrix}\right]\left[\begin{matrix}x \\ y \\z\end{matrix}\right]=\left[\begin{matrix}0\\0 \\0\end{matrix}\right]\]

OpenStudy (phi):

the z (associated with the row of zeros) is a free variable. People usually pick it to have a value of 1 now solve for x and y (I used a and b instead of the long decimals)

OpenStudy (phi):

remember the matrix is short-hand for a system of equations in this case 0*z = 0 (let z=1 , though any value would work, 1 is a nice choice) 0x + 1y + b*z= 0 1 x + 0*y +a*z = 0 solve for y and x

OpenStudy (phi):

you get x= -a, y=-b, z = 1 in other words, the vector is \[\left[\begin{matrix}-a \\ -b \\ 1\end{matrix}\right]\]

OpenStudy (phi):

where in our case, a and b are the entries in the last column of the reduced row echelon matrix

OpenStudy (anonymous):

ok sorry I'm trying to do it on my end .... I'm lost

OpenStudy (phi):

the last step is we want the entries in the vector to sum up to 1 so divide the entries by the sum

OpenStudy (phi):

what step ?

OpenStudy (anonymous):

so x= -1.03544286 y= -.63460141 z= 1

OpenStudy (phi):

almost. the equation (for x for example) is x -1.035*z = 0 where we let z= 1 we get x - 1.035 = 0

OpenStudy (phi):

in other words, we negate the two entries in the 3rd column when we use them in the eigenvector

OpenStudy (anonymous):

ok so how do we divide by the sum

OpenStudy (anonymous):

my assignment is do in a few so I'm trying to understand this quicker

OpenStudy (phi):

add up the entries [ 1.03 0.63 1] (you can use more decimals) then divide each entry by that value

OpenStudy (phi):

and, if you like, multiply that vector by 100, so the values represent percent

OpenStudy (anonymous):

isn't the value 1? if we divide by 1 we will get the same answer??

OpenStudy (phi):

1.03846 0.63462 1.00000 does not add up to 1 but we want it to. so divide each entry by the sum 2.6731

OpenStudy (phi):

it's the same idea as changing [ 1 2] into ⅓ ⅔